-7x+3y=52
15x+6y=-12
Please help! I am supposed to solve this by substitution, and I have no idea how to do it!

Answer by Maths68(1474)

-7x+3y=52................(1)
15x+6y=-12..............(2)
Take (1)
-7x+3y=52................(1)
3y=52+7x
3y/3=(52+7x)/3
y=(52+7x)/3
Put the above value of y in (2)
15x+6y=-12
15x+6((52+7x)/3)=-12
15x+2((52+7x)=-12
15x+104+14x=-12
29x=-12-104
29x=-116
29x/29=-116/29
x=-116/29
Put the value of x in (1)
-7x+3y=52................(1)
-7(-116/29)+3y=52
(812/29)+3y=52
3y=52-(812/29)
3y=(1508-812)/29
3y=696/29
3y=24
3y/3=24/3
y=8
x,y = -116/29, 8
Check
put the values of x and y in (1)
-7x+3y=52................(1)
-7(-116/29)+3(8)=52
812/29+24=52
(812+696)/29=52
1508/29=52
1508=52*29
1508=1508